a) asi d’ rsonval meter movement with an internal resistan e Rm = 100 Ω and a full s ale

current of Im = 1 mA is to be converted into a multiple-range d.c voltmeter with ranges of

0 – 10 V, 0 – 50V, 0 – 250 V and 0 – 500 V. Find the values of the various resistances using

the potential divider arrangement.


1
Expert's answer
2021-08-11T09:54:18-0400

Given that: Internal resistance of the meter "R_{m} = 100\space \Omega"


Full scale deflection of current IFSD "= I_{m} = 1\space mA"


Here, we know that the lowest range of the voltmeter is 0–10 V. Therefore, this range must correspond to the resistance nearest to the meter,


The total resistance "R_{T}" for this voltage range can be obtained as:

"R_T=\dfrac{V}{I_{FSD}}=\dfrac{10}{1\times10^{-3}}=10\space K\Omega"

"R _4\n\u200b\n =R_T\n\u200b\n \u2212R _m\n\u200b\n =10000\u03a9\u2212100\u03a9=9900\u03a9"


Similarly, for 0–50 V range, the total resistance "R_{T}" becomes:

"R_T=\dfrac{V}{I_{FSD}}=\dfrac{50}{1\times10^{-3}}=50\space K\Omega"


Similarly, for 0–250 V range, the total resistance "R_{T}" becomes:

"R_T=\dfrac{V}{I_{FSD}}=\dfrac{250}{1\times10^{-3}}=250\space K\Omega"


Similarly, for 0–500 V range, the total resistance "R_{T}" becomes:

"R_T=\dfrac{V}{I_{FSD}}=\dfrac{500}{1\times10^{-3}}=500\space K\Omega"